# thanks: Current affiliation: ISIS Neutron and Muon Source, STFC Rutherford Appleton Laboratory, Didcot OX11 0QX, UK

### v1

@@ IV Prismatic Analysis and Data Processing
While the above approach works very well, it is interesting to compare it to an alternative method of data analysis which we dub the Positionally-Calibrated Prismatic Analysis (PCPA) technique. As opposed to defining discrete sub-regions in detector space and calibrating the average energy being detected, the alternative approach defines discrete steps for the final neutron energies $E_{f}$ and reconstructs the spatial distribution of each $E_{f}$ on the detector system. This approach is demonstrated in Fig. 5. The rationale for this approach is that neutron detection along the detector is very well-defined by the detector electronics, which for the CAMEA implementation at PSI has 1024 pixels over an active length of $\approx 0.9$ m. Just as with ECPA, calibration with a vanadium standard is essential.
@@
To calibrate the analyzer-detector system, the primary spectrometer is tuned to a particular incident energy $E_{i}$ to analyze the elastic incoherent scattering from the vanadium kernel. We have verified that the $2\theta$-position of the secondary spectrometer does not affect the results; thus, the calibration procedure is performed at a single position of the detector system. After scattering from the vanadium and a given analyzer blade, neutrons with energy $E_{f}=E_{i}$ are detected at particular locations on the PSD. This yields a Gaussian distribution for each $E_{f}$, as seen in Fig. 5. This process is repeated by sweeping $E_{i}$ across various bands of energy centered around the nominal $E_{f}(s)$ of the analyzers, that is $E_{i}=E_{f}(s)\pm 0.12$ meV is scanned in steps of $0.04$ meV for $s=1\cdots N_{s}$ (total of $\approx 7$ $E_{i}$’s per station $s$).
@@
This process provides the distribution of neutron counts versus detector position for each discrete $E_{f}(=E_{i})$ within the calibration set. When comparing the different $E_{f}$’s, it is clear that the total detector-integrated intensities $I(E_{f})$ vary significantly due to the analyzer’s mosaic distribution and the radial arrangement of the detectors, see Fig. 5. To correct for this effect, we introduce a scale factor, $G(E_{f})=I(E_{f}=3.21\,\mathrm{meV})/I(E_{f})$, where $I(E_{f})$ is the integrated intensity of a given $E_{f}$ over the entire detector, therefore summing over all Gaussian profiles ascribed to a specific $E_{f}$. The integrated intensity of the $E_{f}=3.21$ meV energy was arbitrarily chosen as a reference as it is the most intense. The scale factor is then applied as a prefactor to each Gaussian profile in the detector space. The last, and central step is to transpose the ensemble of these scaled distributions (of neutron counts versus position at fixed $E_{f}$) into a new set of distributions representing neutron counts versus $E_{f}$ at fixed position. In other words, we create normalized histograms of final energies for each pixel, representing the likelihood that a neutron detected on a given pixel had energy $E_{f}$. A representative result of such a histogram is shown in Fig. 6. Note that the scaling factor $G(E_{f})$ only affects the relative distribution of energies within a histogram; in the end, the scaling does not affect the net flux at the detector. In the next section, we employ the PCPA technique on a realistic sample kernel for a virtual spectrometer to asses its feasability.
@@ V Results and Secondary Spectrometer Optimization
Figure 7: Scattering intensity represented as constant energy slices for the results of our single-crystal simulations with (a) $E=[0.4,0.6]$ meV, (b) $E=[0.9,1.1]$ meV, and (c) $E=[1.4,1.6]$ meV. These simulations utilize full models for the primary spectrometer and secondary spectrometer as described in Sec. II. Two incident energies, several positions of the detector system, and full rotations of the sample were used to acquire the data which was subsequently analyzed using the Positionally-Calibrated Prismatic Analysis technique described in Sec. IV.
Figure 8: Scattering intensity represented as momentum-energy slices for the results of our single-crystal simulations with the momentum transfer $Q_{x}$ as the scan direction $Q_{\parallel}$ and $Q_{y}$ as $Q_{\perp}$ direction integrated over ±$\pm 0.1$${}^{-1}$ for: (a) $Q_{y}=-2.0$ and (b) $Q_{y}=+0.45$${}^{-1}$. Two incident energies, several positions of the detector system, and full rotations of the sample were used to acquire the data which was subsequently analyzed using the Positionally-Calibrated Prismatic Analysis technique described in Sec. IV.
@@ V.1 Original Model
We now turn to our results of the expected performance of the full MANTA model discussed in Sec. II using the single-crystal kernel and the simulations methods of Sec. III and the PCPA calibration technique discussed in Sec. IV. In Fig. 7, we present the results for a complete set of single-crystal rotations as constant energy-slices representing the neutron scattering intensity in the $(Q_{x},Q_{y},0)$ scattering plane for three different energy transfers spanning the bottom, middle, and top of the dispersion band. In Fig. 8, we present momentum-energy slices across the same datasets. These results highlight the efficiency of the secondary spectrometer in capturing the entire volume of $S(Q_{x},Q_{y},E)$ with only two incident energies $E_{i}$. Furthermore, it reveals resolution effects that are expected on triple-axis spectrometers, namely focusing and de-focusing effects, see e.g. the distribution of intensity on excitation rings in Fig. 7(a) and Fig. 7(b), as well as energy-broadening due to momentum integration and analyzer mosaic, see e.g. Fig 8(b).
@@ V.2 Studies of Analyzer Mosaic Effects
Figure 10: Scattering intensity for different secondary spectrometer designs and a simplified primary spectrometer, represented as momentum-energy slices with momentum transfer $Q_{\parallel}=Q_{x}$ for $Q_{\perp}=Q_{y}=-2.0$. The columns show the effect of increasing analyzer mosaic, starting from a $0.5^{\circ}$ mosaic on the top and ending with a $2^{\circ}$ mosaic at the bottom. Similarly, the rows show the effect of increasing the number of analyzer stations, beginning with 5 analyzer stations on the left and ending with 10 on the right. All designs were tested using 57 values of $2\theta$, 72 angles of $\phi$, and 2 $E_{i}$ values. These results show that the PCPA approach is the most efficient with a coarse crystal mosaic because energies are measured more homogeneously. Additionally, more energies can be analyzed using ${}^{\circ}$ crystal mosaic analyzers, reducing gaps in energy in the data.
@@ V.3 Towards Continuous Energy Coverage
With 8 analyzer stations, there is a gradual increase in the $E_{f}$ spacing of the analyzers, resulting in a non-continuous coverage of energy transfer. This increase in spacing with increasing $E_{f}$ is chosen due to the coarser energy resolution for the larger $E_{f}$ analyzers, thereby minimizing overlap between the energies scattered by differing stations. To eliminate gaps in the measured neutron scattering spectrum, an additional measurement, using, for instance, $E_{i}+0.17$ meV, must be combined with the original $E_{i}$, which doubles the measurement time. However, with the PCPA method, it is, in principle, ideal to include overlap in the energies an analyzer will measure. The calibration accounts for, and in fact improves if multiple analyzers scatter neutrons with the same energy $E_{f}$. This is because the uncertainty of using the scale factor $G(E_{f})$ for normalization is reduced.
@@
Using this guiding principle, we modified the virtual secondary spectrometer to include 10 analyzer stations, using constant steps in the analyzer’s energy $E_{f}$, beginning at $E_{f}=$3.20 meV and increasing in intervals of 0.20 meV up to $E_{f}=5.0$ meV. The upper limit of $E_{f}=5$ meV is constrained by the need to use a beryllium filter to remove higher-order neutron wavelength contamination from the scattered beam, while the lower limit of $E_{f}=3.20$ meV was chosen due to increasingly large scattering angles required for low-energy measurements. To explore a cost-effective alternative, we also created a 5 stations design with even intervals of 0.40 meV over the same $E_{f}$ range. By performing simulations with $E_{i}$ and $E_{i}+0.20$ meV, it is possible to obtain a complete neutron scattering spectrum using the PCPA method.
@@
To compare all three designs at different analyzer mosaic values and reduce computation time, we conducted a similar simulation procedure described in Section III but with a few modifications. First, an idealized, nearly monochromatic virtual source within McStas and a curved monochromator created a beam directed at the vanadium kernel for calibration and simulation of the single crystal kernel. This simplified primary spectrometer was less computationally costly than a full simulation. We repeated the vanadium calibration process in a nearly identical manner for each design. However, for large analyzer mosaics, final energies that were previously negligible need to be accounted for, and we used additional calibration energies to properly interpolate the signal. Finally, all designs used a second $E_{i}$ in addition to $E_{i}=5.1\,\mathrm{meV}$. The 5 station design used $E_{i}=5.3\,\mathrm{meV}$, the 8 station design again used $E_{i}=5.27\,\mathrm{meV}$, and the 10 station design used $E_{i}=5.2\,\mathrm{meV}$. The results, shown in Fig. 10 for a typical momentum-energy slice, show that the integrated count rate increases with mosaic with little visible degradation in resolution. Note, however, that these simulations do not consider the decrease in peak reflectivity, and therefore the increase in total count rate will be partially mitigated. As mentioned previously, the increase in mosaic is not a fully independent parameter and will result in other effects not fully modeled in McStas, such as background from thermal diffuse scattering.
@@
To better quantify the quality of the results, we investigated the single crystal dispersion signal’s energy broadening due to MANTA’s finite energy resolution. We did this by fitting the data in Figs. 8(d) and Fig. 10 to constant-${Q}$ cuts separated by intervals of 0.1 Å$\,\mathrm{\AA^{-1}}$ and subsequently fitted with a Gaussian energy profile using $E$ bins with a width of 0.04 meV. The subsequent fitted Gaussian FWHMs as a function of $Q_{x}$ are shown in Fig. 11. The effective energy broadening on the dispersive signal ranges from 0.2 to 0.3 meV, somewhat coarser than for a cold triple-axis spectrometer, an effect we attribute to our virtual experiments’ open, uncollimated geometry and the integration over a sloped, dispersing signal. Additionally, Fig. 11 shows that the energy broadening between the $1^{\circ}$ mosaic simplified model and the $2^{\circ}$ mosaic simplified model is nearly identical, but taking into account the full primary spectrometer coarsens the energy resolution to up to 0.4 meV. Using prismatic analysis, the dispersion energy resolution appears independent of crystal mosaic, which agrees with the results reported in the original prismatic concept paper [21]. It is also worth noting that despite the potential background challenges from coarse mosaicity, using a mosaic as high as $2^{\circ}$ is not without precedent. The indirect time-of-flight spectrometer FARO [22], which also uses prismatic analyzers, decided to use this coarse mosaic value.
@@
Figure 11: Effective energy resolution measured on a dispersive signal extracted from Figs. 8(b) and 10 after fitting constant-$Q$ cuts in intervals of $Q_{x}=$ 0.1 Å$\,\mathrm{\AA^{-1}}$. The resolution is measured as the FWHM of the resulting Gaussian profile.

### v2

@@ IV Prismatic Analysis and Data Processing
While the above approach works very well, it is interesting to compare it to an alternative method of data analysis which we call the Positionally-Calibrated Prismatic Analysis (PCPA) technique. As opposed to defining discrete sub-regions in detector space and calibrating the average energy being detected, the alternative approach defines discrete steps for the final neutron energies $E_{f}$ and reconstructs the spatial distribution of each $E_{f}$ on the detector system. This approach is demonstrated in Fig. 5. The rationale for this approach is that neutron detection along the detector is very well-defined by the detector electronics, which for the CAMEA implementation at PSI has 1024 pixels over an active length of $\approx 0.9$ m [12]. Just as with ECPA, calibration with a vanadium standard is essential.
@@
To calibrate the analyzer-detector system, the primary spectrometer is tuned to a particular incident energy ($E_{i}$) to analyze the elastic incoherent scattering from the vanadium kernel. It is important to note that the vanadium calibration procedure must use a vanadium sample of the same height as the sample to be measured to properly account for the correct vertical divergence in the beam after scattering. We have verified that the $2\theta$-position of the secondary spectrometer does not affect the results; thus, the calibration procedure is performed at a single position of the detector system. After scattering from the vanadium and a given analyzer blade, neutrons with energy $E_{f}=E_{i}$ are detected at particular locations on the PSD. This yields a Gaussian distribution for each $E_{f}$, as seen in Fig. 5. This process is repeated by sweeping $E_{i}$ across various bands of energy centered around the nominal $E_{f}(s)$ of the analyzers, that is $E_{i}=E_{f}(s)\pm 0.12$ meV is scanned in steps of $0.04$ meV for $s=1\cdots N_{s}$ (total of $\approx 7$ $E_{i}$’s per station $s$).
@@
This process provides the distribution of neutron counts versus detector position for each discrete $E_{f}=E_{i}$ within the calibration set. When comparing the different $E_{f}$’s, it is clear that the total detector-integrated intensities $I(E_{f})$ vary significantly due to the radial arrangement of the detectors and the fact that the additional energies measured from the prismatic concept are intrinsically less intense, see Fig. 5. To correct for this effect, we introduce a scale factor, $G(E_{f})=I(E_{f}=3.21\,\mathrm{meV})/I(E_{f})$, where $I(E_{f})$ is the integrated intensity of a given $E_{f}$ over the entire detector, therefore summing over all Gaussian profiles ascribed to a specific $E_{f}$. The integrated intensity of the $E_{f}=3.21$ meV energy was arbitrarily chosen as a reference as it is the most intense. The scale factor is then applied as a prefactor to each Gaussian profile in the detector space. The last, and central step is to transpose the ensemble of these scaled distributions (of neutron counts versus position at fixed $E_{f}$) into a new set of distributions representing neutron counts versus $E_{f}$ at fixed position. In other words, we create normalized histograms of final energies for each pixel, representing the likelihood that a neutron detected on a given pixel had energy $E_{f}$. A representative result of such a histogram is shown in Fig. 6. Note that the scaling factor $G(E_{f})$ only affects the relative distribution of energies within a histogram; in the end, the scaling does not affect the net flux at the detector. Additionally, to mimic the effect of cross-talk baffles, neutrons that land in between strips of neutrons (as seen in Figure 3) are discarded. In the next section, we employ the PCPA technique on a realistic sample kernel for a virtual spectrometer to assess its feasibility.
@@ V Results and Secondary Spectrometer Optimization
Figure 7: Scattering intensity represented as constant energy slices for the results of our single-crystal simulations with (a) $E=0.5\pm 0.1$ meV, (b) $E=1.0\pm 0.1$ meV, and (c) $E=1.5\pm 0.1$ meV. These simulations utilize full models for the primary spectrometer and secondary spectrometer as described in Sec. II. Two incident energies, several positions of the detector system, and full rotations of the sample were used to acquire the data which was subsequently analyzed using the Positionally-Calibrated Prismatic Analysis technique described in Sec. IV.
Figure 8: Scattering intensity represented as momentum-energy slices for the results of our single-crystal simulations with the momentum transfer $Q_{x}$ as the scan direction $Q_{\parallel}$ and $Q_{y}$ as $Q_{\perp}$ direction integrated over ±$\pm 0.1$-1 for: (a) $Q_{y}=-2.0$ and (b) $Q_{y}=+0.45$-1. Two incident energies, several positions of the detector system, and full rotations of the sample were used to acquire the data which was subsequently analyzed using the Positionally-Calibrated Prismatic Analysis technique described in Sec. IV.
@@ V.1 Original Model
We now turn to our results of the expected performance of the full MANTA model discussed in Sec. II using the single-crystal kernel and the simulations methods of Sec. III and the PCPA calibration technique discussed in Sec. IV. In Fig. 7, we present the results for a complete set of single-crystal rotations as constant energy-slices representing the neutron scattering intensity in the $(Q_{x},Q_{y},0)$ scattering plane for three different energy transfers spanning the bottom, middle, and top of the dispersion band. In Fig. 8, we present momentum-energy slices across the same datasets. These results highlight the efficiency of the secondary spectrometer in capturing the entire volume of $S(Q_{x},Q_{y},E)$ with only two incident energies $E_{i}$. Furthermore, it reveals resolution effects that are expected on triple-axis spectrometers[3], namely focusing and de-focusing effects, see e.g. the distribution of intensity on excitation rings in Fig. 7(a) and Fig. 7(b), as well as energy-broadening due to momentum integration and analyzer mosaic, see e.g. the width of the dispersion (cyan) in Fig 8. The width is calculated in intervals of $\Delta Q_{x}=0.2\,\mathrm{\AA^{-1}}$ as the full-width-half-maximum (FWHM) of the intensity integrated over a width of $Q_{x}\pm 0.05$.
@@ V.2 Studies of Analyzer Mosaic Effects
Figure 10: Scattering intensity for different secondary spectrometer designs and a simplified primary spectrometer, represented as momentum-energy slices with momentum transfer $Q_{\parallel}=Q_{x}$ for $Q_{\perp}=Q_{y}=-2.0$. The columns show the effect of increasing analyzer mosaic, starting from a $0.5^{\circ}$ mosaic on the top and ending with a $2^{\circ}$ mosaic at the bottom. Similarly, the rows show the effect of increasing the number of analyzer stations, beginning with 5 analyzer stations on the left and ending with 10 on the right. All designs were tested using 57 values of $2\theta$, 72 angles of $\phi$, and 2 $E_{i}$ values. Additionally, the width of the dispersion (cyan), calculated in the same manner as in Section 8 is plotted on top of the dispersion. These results show that the PCPA approach is the most efficient with a coarse crystal mosaic because energies are measured more homogeneously. Additionally, more energies can be analyzed using 2 crystal mosaic analyzers, reducing gaps in energy in the data.
@@ V.3 Towards Continuous Energy Coverage
With 8 analyzer stations, there is a gradual increase in the $E_{f}$ spacing of the analyzers, resulting in a non-continuous coverage of energy transfer. This increase in spacing with increasing $E_{f}$ is chosen due to the coarser energy resolution for the larger $E_{f}$ analyzers, thereby minimizing overlap between the energies scattered by differing stations. To eliminate gaps in the measured neutron scattering spectrum, an additional measurement, using, for instance, $E_{i}+0.17$ meV, must be combined with the original $E_{i}$, which doubles the measurement time. However, with the PCPA method, it is, in principle, ideal to include overlap in the energies an analyzer will measure. The calibration accounts for, and in fact improves if multiple analyzers scatter neutrons with the same energy $E_{f}$. This is because the uncertainty from using the scale factor $G(E_{f})$ for normalization is reduced.
@@
Using this guiding principle, we modified the virtual secondary spectrometer to include 10 analyzer stations, using constant steps in the analyzer’s energy $E_{f}$, beginning at $E_{f}=$3.20 meV and increasing in intervals of 0.20 meV up to $E_{f}=5.0$ meV. The upper limit of $E_{f}=5$ meV is constrained by the need to use a beryllium filter to remove higher-order neutron wavelength contamination from the scattered beam, while the lower limit of $E_{f}=3.20$ meV was chosen due to increasingly large scattering angles required for low-energy measurements. To explore a lower cost alternative, we also created a 5 stations design with even intervals of 0.40 meV over the same $E_{f}$ range. By performing simulations with $E_{i}$ and $E_{i}+0.20$ meV, it is possible to obtain a complete neutron scattering spectrum using the PCPA method.
@@
To compare all three designs at different analyzer mosaic values, we employed a “toy model” approach. The toy model follows a similar simulation procedure as described in Section III but with a few modifications to ease computational intensity. First, a planar, circular source with radius of 1 cm was focused towards a doubly focusing monochromator identical to the one planned to be used in the front end [19]. The source-to-monochromator and monochromator-to-sample distance was set to the exact same distance as the planned primary spectrometer, 1.6 meters. As a result, the spot size at the sample location using this simplified front end was nearly identical to the one produced by the full primary spectrometer. All of the designs required separate vanadium calibrations conducted in the same manner as described in Section III, except for a few modifications. First, the source was chosen to be monochromatic to have a more precise calibration for a given $E_{f}$. Additionally, for large analyzer mosaics, final energies that were previously negligible need to be accounted for, thus we used additional calibration energies as needed to properly interpolate the signal. While the simulations employing the vanadium sample for the calibration used an entirely monochromatic source term, the source used to produce the dispersion curve using the inelastic scattering kernel used a bandwidth of energies where ($\frac{\Delta E_{i}}{E_{i}}=0.2$), which produced a similar energy resolution at the sample location as the full primary spectrometer. Finally, all designs used a second $E_{i}$ in addition to $E_{i}=5.1\,\mathrm{meV}$. The 5 station design used $E_{i}=5.3\,\mathrm{meV}$, the 8 station design again used $E_{i}=5.27\,\mathrm{meV}$, and the 10 station design used $E_{i}=5.2\,\mathrm{meV}$. The results, shown in Fig. 10 for a typical momentum-energy slice, show that, in general, the integrated count rate increases with mosaic with little visible degradation in resolution. However, the 10 station $2^{\circ}$ mosaic analyzer design produces artifacts outside the signal dispersion. This may occur from the increased likelihood of scattering from multiple analyzers due to the proximity of the analyzers in the 10 station design and the large acceptance of the analyzer from the use of $2^{\circ}$. Additionally, note that these simulations do not consider the decrease in peak reflectivity, and therefore the increase in total count rate will be partially mitigated. As mentioned previously, the increase in mosaic is not a fully independent parameter and will result in other effects not fully modeled in McStas, such as background from thermal diffuse scattering.
@@
To better quantify the quality of the results, we investigated the single crystal dispersion signal’s energy broadening due to MANTA’s finite energy resolution. We did this by fitting the data in Figs. 8(a) and Fig. 10 to constant-${Q}$ cuts separated by intervals of 0.1 Å$\,\mathrm{\AA^{-1}}$ and subsequently fitted with a Gaussian energy profile using $E$ bins with a width of 0.04 meV. The subsequent fitted Gaussian FWHMs as a function of $Q_{x}$ are shown in Fig. 11. The effective energy broadening on the dispersive signal using the full MANTA spectrometer ranges from 0.2 to 0.35 meV, comparable to the resolution of the CTAX spectrometer [32] and CAMEA [13]. We also see that the energy broadening of the $1^{\circ}$ mosaic toy model and the full spectrometer are nearly identical, confirming the validity of the toy model approach. We also note that apart from the region near $Q_{x}=0$, the $2^{\circ}$ mosaic toy model produces identical energy broadening as the other three designs. Thus, using prismatic analysis, the dispersion energy resolution appears independent of crystal mosaic, which agrees with the results reported in the original prismatic concept paper [21]. It is also worth noting that despite the potential background challenges from coarse mosaicity, using a mosaic as high as $2^{\circ}$ is not without precedent. The indirect time-of-flight spectrometer FARO [22], which also uses prismatic analyzers, decided to use this coarse mosaic value.
While the energy-broadening of the single crystal dispersion signal successfully depicts instrument-induced resolution effects, the $Q$-broadening of the single crystal are dominated by the sloped dispersion signal. The integration volume necessary to collect enough statistics ($\Delta Q_{\parallel}\approx 0.1\,\mathrm{\AA^{-1}}$) is large enough where the dispersing signal is the largest contribution of the calculated full-width-half-maximum (FWHM). For this reason, an elastic Bragg scattering sample was used to calculate the elastic $Q$ broadening using the full primary and secondary spectrometer, as described in Section III. The intensity of the Bragg peak was integrated over a volume of $\Delta E=0.04$ meV and $\Delta Q=0.01\,\mathrm{\AA^{-1}}$, and then the remaining $Q$ resolution was fit to a Gaussian and is tabulated in Tab. 3. The broadening of both $Q_{\parallel}$ and $Q_{\perp}$ (via $\theta$$2\theta$ and transverse scans respectively) were calculated. Additionally, a single slit with an opening of $1^{\circ}$ was placed in front of the analyzer channel and compared to the setup with no slit used in the other simulations.
@@
Figure 11: Effective energy broadening measured on a dispersive signal extracted from Figs. 8(a) and 10 after fitting constant-$Q$ cuts in intervals of $Q_{x}=$ 0.1 Å$\,\mathrm{\AA^{-1}}$. The broadening is measured as the FWHM of the resulting Gaussian profile.
Elastic $Q$ broadening
$\Delta Q_{\perp}$ (Å$(\,\mathrm{\AA^{-1}})$ $\Delta Q_{\parallel}$ (Å$(\,\mathrm{\AA^{-1}})$
$E_{i}=3.21$ meV 0.069 0.080
$E_{i}=3.21$ meV $1^{\circ}$ slit 0.048 0.060
$E_{i}=4.05$ meV 0.051 0.075
$E_{i}=4.05$ meV $1^{\circ}$ slit 0.034 0.045
$E_{i}=5.01$ meV 0.050 0.076
$E_{i}=5.01$ meV $1^{\circ}$ slit 0.030 0.061
Table 3: Elastic $Q$ broadening calculated from an elastic Bragg scattering sample. The $Q$ broadening was calculated by integrating over a volume of $\Delta E=0.04$ meV and $\Delta Q=0.01~{}\,\mathrm{\AA^{-1}}$ in the orthogonal direction of the component of the $Q$ broadening being calculated. We calculated the broadening for both $Q_{\perp}$ and $Q_{\parallel}$ across multiple $E_{i}$ as well as in an open set up and one with a single slit with an opening of $1^{\circ}$.